if the distance of the points P(x,y) from A(a,0) be a+x prove that y2 = 4ax

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
Using distance formula
\large \sqrt[(x-a)² + y²] = x +a
Squaring both the sides
 
(x-a)² + y² = (x +a)²
x² + a² - 2a*x + y² = x² + a² + 2a*x
y² = 4a*x
Hence proved
 
Regards
Arun (askIITians forum expert)